हिंदी
तमिलनाडु बोर्ड ऑफ सेकेंडरी एज्युकेशनएचएससी विज्ञान कक्षा ११

Suppose the chances of hitting a target by a person X is 3 times in 4 shots, by Y is 4 times in 5 shots, and by Z is 2 times in 3 shots. They fire simultaneously exactly one time

Advertisements
Advertisements

प्रश्न

Suppose the chances of hitting a target by a person X is 3 times in 4 shots, by Y is 4 times in 5 shots, and by Z is 2 times in 3 shots. They fire simultaneously exactly one time. What is the probability that the target is damaged by exactly 2 hits?

योग
Advertisements

उत्तर

Given

Probability X hitting the target P(X) = `3/4`

Probability Y hitting the target P(Y) = `4/5`

Probability Z hitting the target P(Z) = `2/3`

`"P"(bar"X")` = 1 – P(X) = `1 - 3/4 = 1/4`

`"P"(bar"Y")` = 1 – P(Y) = `1 - 4/5 = 1/5`

`"P"(bar"Z")` = 1 – P(Z) = `1 - 2/3 = 1/3`

Probability hitting the target exactly by 2 hits

= `"P"[("X" ∩ "Y" ∩ bar"Z") ∪ (bar"X" ∩ "Y" ∩ "Z") ∪ ("X" ∩bar"Y" ∩ "Z")]`

= `"P"("X" ∩ "Y" ∩ bar"Z") + "P"(bar"X" ∩ "Y" ∩ "Z") + "P"("X" ∩bar"Y" ∩ "Z")`

= `"P"("X") "P"("Y") "P"(bar"Z") + "P"(bar"X") "P"("Y") "P"("Z") + "P"("X") "P"(bar"Y") "P"("Z")`

= `3/4 xx 4/5 xx 1/3 + 1/4 xx 4/5 xx 2/3 + 3/4 xx 1/5 xx 2/3`

= `(12 + 8 + 6)/60`

= `26/60`

= `13/30`

shaalaa.com
  क्या इस प्रश्न या उत्तर में कोई त्रुटि है?
अध्याय 12: Introduction to probability theory - Exercise 12.3 [पृष्ठ २५९]

APPEARS IN

सामाचीर कलवी Mathematics - Volume 1 and 2 [English] Class 11 TN Board
अध्याय 12 Introduction to probability theory
Exercise 12.3 | Q 12 | पृष्ठ २५९

संबंधित प्रश्न

A fair coin is tossed five times. Find the probability that it shows exactly three times head.


The probability that a certain kind of component will survive a check test is 0.6. Find the probability that exactly 2 of the next 4 tested components survive


In a game, a man wins Rs 5 for getting a number greater than 4 and loses Rs 1 otherwise, when a fair die is thrown. The man decided to thrown a die thrice but to quit as and when he gets a number greater than 4. Find the expected value of the amount he wins/loses


Determine P(E|F).

Two coins are tossed once, where 

E: no tail appears, F: no head appears


Determine P(E|F).

Mother, father and son line up at random for a family picture

E: son on one end, F: father in middle


Consider the experiment of throwing a die, if a multiple of 3 comes up, throw the die again and if any other number comes, toss a coin. Find the conditional probability of the event ‘the coin shows a tail’, given that ‘at least one die shows a 3’.


Two balls are drawn at random with replacement from a box containing 10 black and 8 red balls. Find the probability that

  1. both balls are red.
  2. first ball is black and second is red.
  3. one of them is black and other is red.

Three cards are drawn at random (without replacement) from a well-shuffled pack of 52 playing cards. Find the probability distribution of the number of red cards. Hence, find the mean of the distribution.


 Two balls are drawn from an urn containing 3 white, 5 red and 2 black balls, one by one without replacement. What is the probability that at least one ball is red?


Two balls are drawn from an urn containing 5 green, 3 blue, and 7 yellow balls one by one without replacement. What is the probability that at least one ball is blue?


From a pack of well-shuffled cards, two cards are drawn at random. Find the probability that both the cards are diamonds when first card drawn is kept aside


Two cards are drawn one after the other from a pack of 52 cards without replacement. What is the probability that both the cards drawn are face cards?


If for two events A and B, P(A) = `3/4`, P(B) = `2/5`  and A ∪ B = S (sample space), find the conditional probability P(A/B)


Two thirds of students in a class are boys and rest girls. It is known that the probability of a girl getting a first grade is 0.85 and that of boys is 0.70. Find the probability that a student chosen at random will get first grade marks.


The total number of ways in which 5 balls of different colours can be distributed among 3 persons so that each person gets at least one ball is ______ 


Two cards are drawn out randomly from a pack of 52 cards one after the other, without replacement. The probability of first card being a king and second card not being a king is:


If for any two events A and B, P(A) = `4/5` and P(A ∩ B) = `7/10`, then `P(B/A)` is equal to ______.


Compute P(A|B), if P(B) = 0.5 and P (A ∩ B) = 0.32.


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×